A watervpump lifts water from a level 10 m below the ground. Water is pumped at a rate of 30kg/minute with negligible velocity and friction. Calculate the minimum horsepower the engine should have for this? Use g=9.8 m/s^2


1
Expert's answer
2022-03-22T15:48:22-0400

Given :-Mass, pump Flow rate = 30 kg / min. = 30 kg / 60 s. = 0.5 kg/s.Height h = 10 m.
we know, pump water has negligible velocity , then it's K.E is zero.Pump work to increase the level of water above 10 m.
Now weight of the water = mg= 0.5 × 9.8 = 4.9 N.Now power = mghP= 4.9 × 10 mP= 49 W.
Now we know 1 hp = 746 W.So minimum power = 49 / 746 =0.066= 6.6 × 10⁻² hp.Hence the minimum horsepower is 6.6 x 10⁻² hp.